# Adaptive Parameter-Free Robust Learning using Latent Bernoulli Variables

### v1

@@ 4 Experiments
The proposed RLVI method is compared to existing approaches in three problem settings: standard parameter estimation, online learning, and deep learning.
@@ 4.1 Benchmark on standard learning problems
Problem
dimension
$n$ $\varepsilon$ $\widetilde{\varepsilon}$
LinReg
${\bm{x}\in\mathbb{R}^{10},y\in\mathbb{R}},\newline \bm{\theta}\in\mathbb{R}^{10}$
40 0.2 0.4
LogReg
${\bm{x}\in\mathbb{R}^{2},y\in\{0,1\}},\newline \bm{\theta}\in\mathbb{R}^{3}$
100 0.05 0.3
PCA
$\bm{z}\in\mathbb{R}^{2},\;\newline \bm{\theta}\in\{\mathbb{R}^{2}:\lVert\bm{\theta}\rVert=1\}$
40 0.2 0.4
Cov
$\bm{z}\in\mathbb{R}^{2},\;\bm{\theta}=(\bm{\mu},\bm{\Sigma})$
50 0.2 0.3
Table 1: Settings for experiments, reproduced from [18]$\bm{\theta}$ is a parametric model; $\bm{x}$, $y$, and $\bm{z}$ denote synthetic data ($n$ samples) contaminated with $\varepsilon\cdot n$ outliers, and $\widetilde{\varepsilon}\geq{\varepsilon}$ is the upper-bound used for Huber, Sever, and RRM.
Figure 2: Linear regression. Left. Average relative error versus varying $\varepsilon$. Right. Box plot of relative error for $\varepsilon=20\%$. Each box spans the $25^{th}$ to $75^{th}$ quantiles; red dots depict the means. For both plots, 100 Monte Carlo runs are used. RLVI obtains higher average accuracy and tighter confidence intervals.
@@
First, we demonstrate that our method is applicable to different maximum likelihood problems and achieves higher model accuracy over contemporary alternatives. Thereto, we reproduce the experiments from [18] comparing various algorithms for robust learning on four problems: linear regression, logistic regression, dimensionality reduction using principal component analysis (PCA), and covariance estimation. Since each problem can be formulated as likelihood maximization, we define the corresponding loss functions as the negative log-likelihood. In the experiments, synthetic data ($n$ samples) is generated from the mixture of $p(\bm{z})$ and $q(\bm{z})$ with a fixed ratio of outliers $\varepsilon$ using the specified model $\bm{\theta}^{\star}$. The optimal model $\widehat{\bm{\theta}}$ is estimated using different algorithms: standard ML, Huber (method tailored for linear regression) [25], Sever [5], RRM [18], and RLVI. The general settings for each problem are summarized in Table 1; the reader is referred to [18] for more details, including specifics of distributions used as true and corrupted. Subsequently, we perform 100 Monte Carlo runs and plot the statistics for corresponding errors with boxplots. For linear regression, the results also include average relative error for $\varepsilon$ varying in $[0;0.4]$ – again, 100 Monte Carlo runs were used for each fixed $\varepsilon$. Results are shown in Figures 2 and 3.
Figure 3: Left. Logistic regression: box plot of angle in degrees between the true separating hyperplane $\bm{\theta}^{\star}$ and estimates $\widehat{\bm{\theta}}$. Middle. PCA: box plot of missalignment errors $1-|\cos{(\widehat{\bm{\theta}}\,^{\top}\bm{\theta}^{\star})}|$ for the subspace spanned by the first principal component. Right. Covariance estimation: box plot of relative errors ${\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{F}/\|\bm{\theta}^{\star}\|_{F}}$, where $\bm{\theta}^{\star}=(\bm{\mu},\bm{\Sigma})$ is the true mean and covariance, and $\widehat{\bm{\theta}}$ is the corresponding estimate.
RLVI achieves better average accuracy and tighter confidence intervals than competing methods. Figure 2 depicts the results of linear regression for different corruption levels. It can be seen that RLVI performs slightly but consistently better than other approaches, including the case of no corruption $\varepsilon=0$, owing to the ability of RLVI to automatically adjust to an arbitrary number of corrupted samples.
@@ 4.2 Online learning
We consider the Human Activity Recognition dataset from [8] containing $24,075$ measurements from smartphone sensors. Each measurement $\bm{z}_{i}=(\bm{x}_{i},y_{i})$ consists of $\bm{x}_{i}\in\mathbb{R}^{60}$ features extracted from accelerometers during different human activities: Sitting, Standing, Walking, Running, and Dancing. We perform binary classification and towards that end, partition the five initial labels into two: resting state ($y=0$) and active state ($y=1$). To simulate a data stream, at each iteration we retrieve data in batches of size $2b$, where $b=100$ samples are used for training and another $b=100$ samples serve performance evaluation. To introduce noise, we corrupt the data in each training batch by randomly flipping $\varepsilon$ percent of positive labels. Moreover, the corruption level $\varepsilon$ changes in each iteration. Since we are testing the robustness of RLVI against corruption, we focus specifically on variation of noise – not on the gradual change of $\bm{\theta}$ (concept drift) or other possible challenges related more to the method to learn $\bm{\theta}$ incrementally. For each batch, $\varepsilon$ is sampled from a linearly transformed Beta distribution (called PERT [12]) that is simple to parameterize with three values, so that the samples are within the interval $[\varepsilon_{\text{min}},\varepsilon_{\text{max}}]$ with the mode $\varepsilon_{\text{mode}}$. To simulate a typical case, we set ${\varepsilon_{\text{min}}=0}$, ${\varepsilon_{\text{max}}=0.3}$, and ${\varepsilon_{\text{mode}}=0.1}$. Thus, the ratio of corrupted labels in each batch varies according to the distribution in Figure 4.
Figure 4: Online classification. Left. Distribution of corruption level across batches of streaming data. Right. Recall (true positive rate) on left-out data versus total number of observed samples – the metric is smoothed with a moving average filter.
Since RRM is also based on block-wise optimization, it can similarly be implemented as incremental learning, where inference for $\bm{\theta}$ consists of anti-gradient steps minimizing the modified empirical risk for each batch. Hence, the experiments compare three versions of online classification: the standard stochastic maximization of the likelihood (SGD), incremental minimization of the risk function defined in RRM [18] with a threshold for $\varepsilon$ set to ${\varepsilon_{\text{max}}=0.3}$, and the presented variational inference approach (RLVI).
@@ 4.3 Overparameterized model
We now extend RLVI to learning an overparameterized model and, as an example, consider image classification using a convolutional neural network and corrupted training labels.
@@
Existing approaches. State-of-the-art performance is achieved in this setting by methods that identify and distill corrupted samples based on some criterion. The algorithm Co-teaching [7], for example, is based on training two neural networks in parallel and learning only on the samples that attain a small loss for both models. The JoCoR approach [20] also trains two models but aims to reduce the diversity between their predictions. Due to simultaneous training of two models, Co-teaching and JoCoR effectively double the computational time. CDR [22] in contrast, trains one model – it employs weight decay to diminish the impact of network parameters that overfit to corrupted samples, and early-stopping using a validation set. The recent algorithm USDNL [23] estimates prediction uncertainty for training samples using dropout in a single network, thus identifying the samples with corrupted labels. Each of the above methods combines its own criterion for outliers with a schedule to gradually consider fewer samples and thus account for model overfitting occurring in later epochs. This schedule is a non-decreasing function defined with the corruption level $\varepsilon$ (assumed to be known in advance) so that, by the end of training, only $(1-\varepsilon)$ ratio of the training set is being used.
Figure 5: Image classification on CIFAR10 corrupted with synthetic noise (pairflip with $\varepsilon=45\%$). Top. Accuracy on the clean testing set for standard SGD and RLVI. Bottom. Percentage of corrupted and non-corrupted samples correctly identified with the decision rule (${\pi_{i}<\tau}$). In both plots, the dashed line corresponds to RLVI with no regularization – solid line indicates that truncation (${\pi_{i}<\tau\implies\pi_{i}=0}$) is used to set the corresponding terms in anti-gradient updates to zero. Decision boundary $\tau$ is computed from the bounded type II error criterion (20). Regularization based on reducing type II error minimizes overfitting, making differentiation of corrupted and non-corrupted data points more effective, ultimately improving test accuracy.
Regularization. As described in Section 3, to prevent overfitting, we implement regularization using hard truncation: ${\pi_{i}<\tau\implies\pi_{i}=0}$. Consequently, we need to evaluate when to apply this regularization, and how to select the truncation boundary $\tau$. The answer to the first question is straightforward: regularization is to be applied when the model begins to overfit, which in practice can be identified with a drop in prediction accuracy on a noisy validation set. As studied empirically [11, 17, 22], neural networks first fit the true distribution. In other words, before the network reaches an overfitting stage, the loss on corrupted samples is not reduced, and computed quantities $\pi_{i}$ allow for differentiation of outliers as described in the first sections.
To answer the second question, we note that in light of corrupted samples, regularization of an overparameterized model gives rise to asymmetrical errors. On one hand, if a fraction of non-corrupted samples is identified as corrupted and thus is excluded from optimization with the decision rule ${\pi_{i}<\tau\implies\pi_{i}=0}$, the loss is minimized on a smaller subset of ‘clean’ samples. Consequently, the model does not fit the training set perfectly. In contrast, if corrupted samples are treated as ‘clean’, the model adapts to the contamination term $q(\bm{z})$ thereby learning the wrong pattern. To account for a higher cost of the latter, type II error, the optimal threshold $\tau^{\star}$ is obtained from the condition that the error is fixed at the admissible level of 5% (typically a standard choice in statistical test theory). Furthermore, the type II error is approximated using the obtained posterior $r(\bm{t}|\bm{\pi})$ from the variational inference formulation,
$\dfrac{\mathbb{E}_{r}\Big{[}\#\text{False Clean}\Big{]}}{\mathbb{E}_{r}\Big{[}\#\text{Corrupted \Big{]} }}=\dfrac{\sum\limits_{i=1}^{n}(1-\pi_{i})\mathbbm{1}[\pi_{i}\geq\tau]}{\sum\limits_{i=1}^{n}(1-\pi_{i})}\leq 0.05.$ (20)
@@
Here, the numerator equals the expected number of corrupted samples considered as ‘clean’, and the denominator is the expected total number of corrupted samples, with respect to posterior probabilities $\bm{\pi}$. To increase the number of samples to be used for learning, one can select the largest $\tau\in\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$ satisfying the inequality above. Such a solution can be readily found after sorting the array $\bm{\pi}\in[0;1]^{n}$ in $O(n\ln n)$ time. Figure 5 presents an example of image classification on CIFAR10 with randomly flipped labels (read about ‘pairflip’ synthetic noise below). It illustrates how the introduced regularization functions in practice: if no regularization is used, the proportion of identified noisy observations decreases as the model gradually fits $q(\bm{z})$, thus attaining lower test accuracy. But with regularization, differentiation of corrupted and non-corrupted data points according to $\pi_{i}$ remains effective during all iterations, which leads to better generalization. In the Appendix, we provide similar plots for other types and levels of noise.
Resulting algorithm. We now introduce Algorithm 2 that implements RLVI as stochastic gradient optimization of neural network parameters $\bm{\theta}$. It achieves robustness by maximizing the evidence lower-bound (15) instead of regular maximization of the likelihood. The latter employs parameters $\pi_{i}$ that represent probability of each sample to be non-corrupt. These parameters are updated at the end of each epoch using efficient fixed-point iterations (17). To prevent model’s overfitting to corrupted samples, the algorithm eliminates gradient terms corresponding to low $\pi_{i}$. Threshold for truncation is re-computed across epochs as a non-decreasing value from the type II error criterion (20).
@@
4:$L\leftarrow$ empty array of size $n$triangleright to store $\ell_{\bm{\theta}}(\bm{z}_{i}),\bm{z}_{i}\in Z_{tr}$
@@
7:for $\text{epoch}=1,2,\dots,n_{\text{epochs}}$ do
8:     for $\mathcal{I}_{b}\sim U\{1,\dots,n\}$ do $\triangleright$ batch of $b$ indices
9:         $L_{i}\leftarrow\ell_{\bm{\theta}}(\bm{z}_{i}),\;i\in\mathcal{I}_{b}$triangleright store loss value
@@
14:         $\tau\leftarrow\max(\tau,\tau^{\star})$triangleright can only increase
15:         $\pi_{i}\leftarrow\pi_{i}$ if $\pi_{i}\geq\tau$ else 00
16:     else
17:         $overfit\leftarrow$ True if accuracy on $Z_{val}$ dropped
18:Output: $\bm{\theta}$
Synthetic noise. To test the algorithm, we conduct the experiments in image classification using convolutional neural networks and cross-entropy loss as $\ell_{\bm{\theta}}$. The datasets being used are MNIST [14], that contains hand-written digits, and CIFAR10 and CIFAR100 [13] – both containing the same images that are classified into ten and one hundred categories respectively. Subsequently, the training labels in these datasets are corrupted with three types of synthetic noise: symmetric, asymmetric, and pairflip. These noise types have been commonly used in previous works [7, 20, 22, 23]. Training was performed with the conventional values of hyperparameters for each case. Namely, for MNIST, we used the LeNet architecture and the SGD optimizer with momentum 0.9, weight decay of strength $10^{-3}$, and linearly decaying learning rate (initial rate 0.01). LeNet was trained for 100 epochs with the batch size 32. For CIFAR10 and CIFAR100, the architecture used was ResNet18 and ResNet34 respectively, trained with the SGD optimizer, with momentum 0.9 and weight decay of strength $5\cdot 10^{-4}$, that employed cosine annealing to decay a learning rate (initial rate 0.01). Optimization was performed for 200 epochs with batches of size 128. Furthermore, we compare the attained classification accuracy on the test set with the standard likelihood maximization approach and four recently proposed alternative methods: Co-teaching, JoCoR, CDR, and USDNL. For CDR and RLVI, 10% of the training data was left-out as a validation set. For RLVI, we identify overfitting, and thus start applying regularization, as soon as validation accuracy at the current epoch becomes less than the average of its two previous values. In contrast, CDR uses a validation set for early-stopping: optimization should conclude if the validation accuracy exceeds some specified threshold. To avoid any uncertainty in selecting this threshold, for CDR, we report the test accuracy that corresponds to the lowest validation loss during all epochs. For all other methods, we consider accuracy at the last epoch as a final result. For Co-teaching, JoCoR, CDR, and USDNL, we used their default hyperparameters related to robust learning, including the same schedule for the ratio of considered samples during epochs, as defined in [7], deduced from the true noise level employed. Also, since USDNL is based on uncertainty estimation using dropout, for USDNL specifically, we used variations of LeNet, ResNet18, and ResNet34 with dropout layers, where dropout rate was set to 0.25 (as in the original paper [23] – but for another architecture).

### v2

@@ 4 Experiments
The proposed RLVI method is compared to existing approaches in three problem settings: standard parameter estimation, online learning, and deep learning [^1] Implementation of RLVI and our experiments are available at https://github.com/akarakulev/rlvi..
Figure 2: Box plots of relative errors for a fixed value of corruption level $\varepsilon$. Left. Linear regression: relative errors ${\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{2}/\|\bm{\theta}^{\star}\|_{2}}$. Middle. Logistic regression: angle in degrees between the true separating hyperplane $\bm{\theta}^{\star}$ and estimates $\widehat{\bm{\theta}}$. Right. PCA: misalignment errors $1-|\cos{(\widehat{\bm{\theta}}\,^{\top}\bm{\theta}^{\star})}|$ for the subspace spanned by the first principal component. Each box spans the 25th to 75th quantiles; red dots depict the means. For all plots, 100 Monte Carlo runs are used.
@@ 4.1 Benchmark on standard learning problems
First, we demonstrate that our method is applicable to different maximum likelihood problems and achieves higher model accuracy over contemporary alternatives. Thereto, we reproduce the experiments from (Osama et al., 2020) comparing various algorithms for robust learning on three problems: linear regression, logistic regression, and dimensionality reduction using principal component analysis (PCA). Since each problem can be formulated as likelihood maximization, we define the corresponding loss functions as the negative log-likelihood. In the experiments, synthetic data ($n$ samples) is generated from the mixture of $p(\bm{z})$ and $q(\bm{z})$ with a fixed ratio of corrupted samples $\varepsilon$ using the specified model $\bm{\theta}^{\star}$. The optimal model $\widehat{\bm{\theta}}$ is estimated using different algorithms: standard ML, Huber (Zoubir et al., 2018), Sever (Diakonikolas et al., 2019), RRM (Osama et al., 2020), and RLVI. The reader is referred to (Osama et al., 2020) for more details on the three test problems, including specifics of distributions used as true and corrupted. Subsequently, we perform 100 Monte Carlo runs and plot the statistics for corresponding errors with boxplots in Figure 2. For linear regression the results, shown in Figure 3, also include average relative error for $\varepsilon$ varying in $[0;0.4]$. Again, 100 Monte Carlo runs are used for each fixed $\varepsilon$.
Problem dimension $n$ $\varepsilon$ $\widetilde{\varepsilon}$
LinReg $\bm{\theta}\in\mathbb{R}^{10}$ 40 0.2 0.4
LogReg $\bm{\theta}\in\mathbb{R}^{3}$ 100 0.05 0.3
PCA $\bm{\theta}\in\{\mathbb{R}^{2}:\lVert\bm{\theta}\rVert=1\}$ 40 0.2 0.4
Table 1: Experiments reproduced from (Osama et al., 2020): $\bm{\theta}$ is a parametric model learned from synthetic data ($n$ samples in total, $\varepsilon n$ are corrupted); $\widetilde{\varepsilon}\geq{\varepsilon}$ is the upper-bound used for Huber, Sever, and RRM.
Note that using the unbounded likelihood for defining $\ell_{\bm{\theta}}(\bm{z})$ within RLVI can lead to a degenerate solution and thus might require additional regularization – see an example with a covariance estimation problem in the Appendix.
From Figures 2 and 3, one can see that RLVI achieves better average accuracy and tighter confidence intervals than the competing methods.
Figure 3: Linear regression. Average relative error versus varying corruption level $\varepsilon$; 100 Monte Carlo runs are used.
@@ 4.2 Online learning
We consider the Human Activity Recognition dataset from (Helou, 2023) containing $24,075$ measurements from smartphone sensors. Each measurement $\bm{z}_{i}=(\bm{x}_{i},y_{i})$ consists of $\bm{x}_{i}\in\mathbb{R}^{60}$ features extracted from accelerometers during different human activities: Sitting, Standing, Walking, Running, and Dancing. We perform binary classification and towards that end, partition the five initial labels into two: resting state ($y=0$) and active state ($y=1$). To simulate a data stream, at each iteration we retrieve data in batches of size $2b$, where $b=100$ samples are used for training and another $b=100$ samples serve performance evaluation. To introduce noise, we corrupt the data in each training batch by randomly flipping $\varepsilon$ percent of positive labels. Moreover, the corruption level $\varepsilon$ changes in each iteration. Since we are testing the robustness of RLVI against corruption, we focus specifically on variation of noise – not on the gradual change of $\bm{\theta}$ (concept drift) or other possible challenges related more to the method to learn $\bm{\theta}$ incrementally. For each batch, $\varepsilon$ is sampled from a linearly transformed Beta distribution (called PERT (Johnson et al., 1995)) that is simple to parameterize with three values, so that the samples are within the interval $[\varepsilon_{\text{min}},\varepsilon_{\text{max}}]$ with the mode $\varepsilon_{\text{mode}}$. To simulate a typical case, we set ${\varepsilon_{\text{min}}=0}$, ${\varepsilon_{\text{max}}=0.3}$, and ${\varepsilon_{\text{mode}}=0.1}$. Thus, the ratio of corrupted labels in each batch varies according to the distribution in Figure 4.
Figure 4: Online classification. Left. Distribution of corruption level across batches of streaming data. Right. Recall (true positive rate) on left-out data versus total number of observed samples (smoothed with a moving average filter).
Since RRM is also based on block-wise optimization, it can similarly be implemented as incremental learning, where inference for $\bm{\theta}$ consists of anti-gradient steps minimizing the modified empirical risk for each batch. Hence, the experiments compare three versions of online classification: the standard stochastic maximization of the likelihood (SGD), incremental minimization of the risk function defined in RRM (Osama et al., 2020) with a threshold for $\varepsilon$ set to ${\varepsilon_{\text{max}}=0.3}$, and the proposed approach, RLVI.
@@ 4.3 Overparameterized model
We now extend RLVI to learning an overparameterized model and consider image classification using a convolutional neural network when training labels are corrupted.
@@
Existing approaches. State-of-the-art performance is achieved in this setting by methods that identify and distill corrupted samples based on some criterion. The algorithm Co-teaching (Han et al., 2018) is based on training two neural networks in parallel and learning only on the samples that attain a small loss for both models. The JoCoR approach (Wei et al., 2020) also trains two models but aims to reduce the diversity between their predictions. Due to simultaneous training of two models, Co-teaching and JoCoR effectively double the computational time. CDR (Xia et al., 2020) trains one model – it employs weight decay to diminish the impact of network parameters that overfit to corrupted samples, and early-stopping using a validation set. The recent algorithm USDNL (Xu et al., 2023) estimates prediction uncertainty for training samples using dropout, thus identifying the samples with corrupted labels. Each of the above methods combines its own criterion for corrupted samples with a schedule to gradually consider fewer samples and thus account for model overfitting occurring in later epochs. The schedule is a non-decreasing function defined with the corruption level $\varepsilon$ (assumed to be known in advance) so that, by the end of training, only $(1-\varepsilon)$ ratio of the training set is being used. The algorithm BARE from (Patel & Sastry, 2023) aims to robustly train neural networks for classification independently of $\varepsilon$. It removes samples from the loss function using batch statistics, assuming that the class conditional noise has a special structure.
[…]
@@
Figure 5 presents an example of image classification on CIFAR10 with randomly flipped labels. It illustrates how the introduced regularization functions in practice: if no regularization is used, the proportion of identified corrupted observations decreases while model gradually fits $q(\bm{z})$ despite marginal likelihood formulation, thus attaining lower test accuracy. But with regularization, differentiation of corrupted and non-corrupted data points according to $\pi_{i}$ is more effective during all iterations, which leads to better generalization. In the Appendix, we provide similar plots for various types and levels of noise.
Figure 5: Image classification on CIFAR10 corrupted with synthetic noise (pairflip, $\varepsilon=45\%$). Top. Accuracy on the clean testing set for standard SGD and RLVI. Bottom. Percentage of corrupted and non-corrupted samples correctly identified with the decision boundary ${\pi_{i}<\tau}$. In both plots, the dashed line corresponds to RLVI with no regularization – solid line indicates that truncation (${\pi_{i}<\tau\implies\pi_{i}\leftarrow 0}$) is used for the terms in anti-gradient updates. Threshold $\tau$ is computed from (22). Regularization based on the bounded type II error makes differentiation of corrupted samples more effective, ultimately improving test accuracy for the overparameterized setting.
@@
4:  $L\leftarrow$ empty array of size $n$ // to store $\ell_{\bm{\theta}}(\bm{z}_{i}),\bm{z}_{i}\in Z_{tr}$
@@
7:  for $epoch=1,2,\dots,n_{\text{epochs}}$
8:     for $\mathcal{I}_{b}\sim U\{1,\dots,n\}$ // for a batch of $b$ indices
9:      $L_{i}\leftarrow\ell_{\bm{\theta}}(\bm{z}_{i}),\;i\in\mathcal{I}_{b}$ // store loss value
@@
14:      $\tau\leftarrow\max(\tau,\tau^{\star})$ // can only increase
15:      $\pi_{i}\leftarrow\pi_{i}$ if $\pi_{i}\geq\tau$ else 00 // regularization
16:     else if accuracy on $Z_{val}$ dropped
17:      $overfit\leftarrow$ True
18:  end for
19:  Output: $\bm{\theta}$
Synthetic corruption. To test the algorithm, we conduct the experiments in image classification using convolutional neural networks and cross-entropy loss as $\ell_{\bm{\theta}}$. The datasets being used are MNIST (LeCun, 1998), that contains hand-written digits, and CIFAR10 and CIFAR100 (Krizhevsky, 2009) – both containing the same images that are classified into ten and one hundred categories respectively. Subsequently, the training labels in these datasets are corrupted with four types of synthetic noise: symmetric, asymmetric, and pairflip (class-dependent) and instance (feature-dependent). These noise types have been commonly used in previous works (Han et al., 2018; Wei et al., 2020; Xia et al., 2020; Xu et al., 2023). Additionally, we perform the experiment in which data contains both the corrupted labels and the out-of-distribution samples. To this end, we consider the dataset CIFAR80N-O (Yao et al., 2021) which is obtained from CIFAR100 as follows: the last 20 classes in CIFAR100 are regarded as out-of-distribution images, and images from the remaining 80 classes are subsequently corrupted by one of the class-dependent synthetic noise types (pairflip, symmetric, and asymmetric). Thereto, we demonstrate that, owing to the generality of Huber contamination model (2), RLVI can be succesfully applied to learning in the presence of noise of arbitrary structure. In these experiments, we employ commonly used hyperparameter settings found in literature specific to each dataset and model architecture. These settings can be found in Table 4 of the Appendix.
We compare the attained classification accuracy on the test set with the standard likelihood maximization approach and recently proposed alternative methods: Co-teaching, JoCoR, CDR, USDNL, and BARE. For CDR and RLVI, 10% of the training data is used as a validation set: in RLVI, we apply regularization (21) after validation accuracy at the current epoch becomes less than the average of its two previous values. In contrast, CDR uses a validation set for early-stopping: optimization concludes if the validation accuracy exceeds some specified threshold. The latter implies that for CDR we report the test accuracy corresponding to the lowest validation loss. For all the alternative methods, we use their default hyperparameters related to robust learning, including the same schedule for the ratio of considered samples during epochs, as defined in (Han et al., 2018), deduced from the true noise level employed. Also, since USDNL is based on uncertainty estimation using dropout, for USDNL specifically, we used variants of the corresponding neural nets with dropout layers, where dropout rate was set to 0.25, as in the original paper (Xu et al., 2023).